OpenStudy (anonymous):

1/x+1/y=1/z solve for x

OpenStudy (anonymous):

1/x+1/y=1/z y/x+1=1/z x+1/y=z x+1=yz x=yz-1

OpenStudy (anonymous):

the answer given is yz/y-z=x i just dont know how to get there

OpenStudy (anonymous):

read the question carefully again